OpenStudy (anonymous):

PLEASE HELP ME 1. At a fundraiser breakfast, the bill for three glasses of orange juice and five pancake specials is $7.60, whereas the bill for one glass of orange juice and two pancake specials is $2.90. a. Write a system of two equations to model this problem. Let j stand for the price of a glass of juice and let p stand for the price of a pancake special. b. Use the substitution method to solve the system of equations. c. What would be the bill for one glass of orange juice and one pancake special?

OpenStudy (anonymous):

@texaschic101 @SolomonZelman

OpenStudy (danjs):

3j+5p=7.60

OpenStudy (danjs):

j +2p = 2.90

OpenStudy (solomonzelman):

let number of glasses of orange juice = J let number of pancake specials = P So, " the bill for three glasses of orange juice and five pancake specials is $7.60" gives you the equation, `3J+5P=7.6` and " whereas the bill for one glass of orange juice and two pancake specials is $2.90." gives you the equation, ` J+2P=2.9`

OpenStudy (solomonzelman):

So ``` your system is 3J+5P=7.6 J+2P=2.9 :) ```

OpenStudy (anonymous):

A.) 3j+5p=7.6 J+2p=2.9 B.) j=2.9-2p 3(2.9-2p)+5p=7.6 8.7-6p+5p=7.6 7.6-8.7=-p P=1.1 Plug in J+2(1.1)=2.9 J=0.7 C.) p+j=? 1.1+0.7=1.8 $1.80

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thank you @SolomonZelman :)

OpenStudy (solomonzelman):

I was away, helping my relatives. Sorry.